A transformer changes 120 V across the primary to 1200 V across the secondary. If the secondary coil has 1556, how many turns does the primary coil have?
Answer #1

Here, for the voltage

Vs = 1200 V

Vp = 120 V

number of turns , Ns = 1556

Now, for the transformer

Ns/Np = Vs/Vp

1556/Np = 1200/120

Np = 155.6

the number of turns needed in primary is 155.6